OpenStudy (aaronandyson):

Perform the indicated expression: 2tan28°tan62°

OpenStudy (ankit042):

x= 28 2tan(90-x).tan(x)= 2sin(90-x).sin(x)/cos(90-x).cos(x)

OpenStudy (ankit042):

Now sin(90-x)=? cos(90-x)=?

OpenStudy (michele_laino):

we have this: \[\Large \tan 90 = \tan \left( {28 + 62} \right) = \frac{{\tan 28 + \tan 62}}{{1 - \tan 28\tan 62}}\]

OpenStudy (jhannybean):

Ahh, I didn't know you could combine it :)

OpenStudy (michele_laino):

now, since tan90= infinity, we have to request that: \[\Large 1 - \tan 28\tan 62 = 0\]

OpenStudy (michele_laino):

therefore: \[\Large 2 - 2\tan 28\tan 62 = 0\]

OpenStudy (aaronandyson):

thanks

OpenStudy (aaronandyson):

1/sec^2(36) - cot^(54) =

OpenStudy (aaronandyson):

^ help me with that